OpenStudy (lena772):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $80 million after 4 years while profits increased $60 million per year. Select the correct graph of this function. There is a graph of a line that has a y intercept that is approximately negative $160 million and an x intercept that is approximately 2.7 years.

OpenStudy (lena772):

There is a graph of a line that has a y intercept of $0 and an x intercept that is 0 years. The line passes through the point x = 2.5 and y = $50 million. There is a graph of a line that has a y intercept that is approximately negative $40 million and an x intercept that is approximately 1.3 years. There is a graph of a line that has a y intercept that is approximately $120 million and passes through the point x = 2 years and y = $100 million.

ganeshie8 (ganeshie8):

$80 million after 4 years (4, 80) is one point on the graph

ganeshie8 (ganeshie8):

profits increased $60 per year that means slope, m = 60

ganeshie8 (ganeshie8):

you know a point and slope write the equation

OpenStudy (lena772):

y=60x+b 80=60*4+b 80=240+b -160=b

OpenStudy (lena772):

y=60x-160

ganeshie8 (ganeshie8):

so, y intercept, b = -160 x intercept = ?

OpenStudy (lena772):

0=60x-160 160=60x 2 2/3 = x

ganeshie8 (ganeshie8):

thought u prefer decimal numbers ?

ganeshie8 (ganeshie8):

0=60x-160 160=60x 2 2/3 = x 2.7 = x

OpenStudy (lena772):

Lol ! You think I prefer calculator :p

OpenStudy (lena772):

So A is our answer?

ganeshie8 (ganeshie8):

Yup

OpenStudy (solomonzelman):

to connect the 2 http://openstudy.com/study#/updates/53234dc0e4b06a39f860fe21